# Cole: A Hierarchical Generation Framework for Graphic Design

### 28 Nov 2023 v1

@@ 3.4 Text-to-Object Diffusion Model
Subsequently, we propose training a text-to-object diffusion model capable of perceiving both the object captions present in the JSON file and the visual pixels predicted by the text-to-background diffusion models.
@@
Following the instruct-P2P approach [3], we concatenate the noise latent vectors with the latent vectors of the conditioning background image with the alpha mask for the object layer, and we add additional input channels to the first convolutional layer. All available weights of the diffusion model are initialized from pre-trained checkpoints, while weights operating on the newly added input channels are initialized to zero. As a result, we can predict both the object image and its alpha mask layer, which enables us to control the transparency of the object pixels. Consequently, we can seamlessly blend the object image layer with the background image layer in accordance with the alpha mask. Empirically, we found that predicting the additional composed image is pivotal to ensuring a more reasonable placement of the object pixels. Therefore, our text-to-object diffusion model needs to predict channels. For more details, please refer to the experimental section.
@@
Similar to the text-to-background generation model, we optimize our assembled dataset, comprising approximately caption and object image pairs. The relatively smaller number of pairs is due to the frequent absence of the object image layer in many graphic design images. We employ the instruct-P2P scheme for optimization, which is based on the DeepFloyd [19] IF diffusion models, and this process is carried out over an estimated span of epochs. Figure 5 provides a selection of representative examples that demonstrate our method of decomposing complex graphic design images into distinct layers of background and object images. It’s important to note that predicting the object alpha mask is essential for facilitating flexible editing operations on the object pixels. However, ensuring accurate alpha mask prediction presents a complex challenge. Please refer to the supplementary materials for more in-depth technical details.
3.5 Typography LMM
Typography [38] is the art and technique of arranging type to make written language legible, clear, and visually engaging for the reader. More succinctly, typography refers to the style and appearance of text. It encompasses the design of font style, appearance, and structure, with the goal of eliciting specific emotions and conveying certain messages. In essence, typography brings text to life.
@@ 3.5 Typography LMM
Contrary to recent efforts, such as Tanveer et al. [32], Iluz et al. [14], and He et al. [10], which primarily employ diffusion models to generate novel letter-level or word-level semantic or artistic typographies, our focus is more directed towards the utilization of an LMM (Large Multimodal Model). This approach allows us to address the highly challenging task of image-level typography. This task involves the selection of font types, establishment of hierarchy arrangements, adjustment of leadings (line spacing), tracking (character spacing), and kernings (space between specific characters).
@@
In our experiments, we elected to fine-tune the LLaVA--B model [24] using nearly triplets for epochs. These triplets consist of multiple texts, composed images of backgrounds and objects, and typography information. Figure 6 displays a representative example of a Typography JSON file. It is evident that predicting visually appealing typography information for each text box, based on the input image, is a complex task. This is due to the requirement of predicting approximately attributes for each box. To our knowledge, even the most recent effort [30] only manages to predict typography attributes, assuming the other attributes are fixed. To simplify the prediction process, we further categorize the continuous floating numbers into discrete bins as per the method outlined in [5] and we also utilize a unique token to represent the values that fall within the same bin. We explain more details of the bin partition scheme in the supplementary material.
3.6 Graphic Render
As demonstrated in the preceding sections, we have successfully predicted the JSON file, complete with detailed captions and text elements. We have also generated the background and object image layers (with object alpha mask), ensuring they adhere to the principles of design aesthetics and layout. In addition, we have produced the typography JSON file, which precisely outlines the style and appearance of the text elements.
We implement a graphic renderer founded on Skia [31], an open-source 2D graphics library. It is important to highlight that the rendering quality, as per the official implementation of CanvasVAE [43], presents significant potential for enhancement. Figure 7 showcases a qualitative comparison of the results obtained from our rendering implementation, contrasted with those generated by CanvasVAE. Upon evaluation, it is evident that our graphic renderer significantly outperforms the open-sourced CanvasVAE.
Original
Ours
CanvasVAE
Original
Ours
CanvasVAE
@@ 3.7 Reflect LMM & Quality LMM
We find that initial graphic design outputs may not always align with user expectations. Therefore, we propose the integration of an added Reflect-LMM. This component is designed to critically analyze the generated design images and predict necessary adjustments to the JSON file, thereby further improving the resulting graphic designs.
The entire self-reflection framework is illustrated on the left side of Figure 8. As an integral part of our implementation plan, we fine-tune a LLaVA--B model on a dataset that comprises more than artificially generated noisy JSON files. Each of these files is paired with a corresponding ground-truth JSON file. Our primary modification involves the addition of random shift noise to the position of each ground-truth typography block.
Besides, we utilize GPT-4V(ision) [25] to evaluate the quality of the generated graphic design images, as displayed in the “Quality Assurance Prompt for GPT-4V(ision)” in the appendix. Through empirical analysis, we have found that GPT-4V(ision) excels at evaluating the quality of main aspects such as design and layout, content relevance, graphics and images, and innovation. However, it falls short in assessing the quality of typography and color. In addition, we also incorporate the evaluation prompt design as introduced in a recent concurrent effort [18]. This evaluates the graphic design images from approximately aspects, including text rendering quality, composition and layout, color harmony, cinematography, style, image-text alignment, aesthetics, and overall design.
4 Experiment
(a) of text box, words, chars
(b) Wordcloud
@@ 4 Experiment
(a) covers & headers
@@
background image
object image
object alpha mask
composed image
To assess our Cole’s proficiency in generating graphic design images, we first build a DesignerIntention benchmark comprising over approximately graphic design intention prompts. These prompts serve as a tool to evaluate the model’s performance from diverse perspectives. Second, as mentioned in the previous section, we propose utilizing GPT-4V(ision) to evaluate the quality of the generated graphic design images and compare our system with the latest systems including DeepFloyd/IF(+GPT-), SDXL(+GPT-), and DALLE3 (+GPT-) following [2, 18]. Last, we carry out ablation experiments to scrutinize the performance of each component and discuss the limitations of our system.
4.1 DesignerIntention Benchmark
We have constructed a benchmark for designer intentions, amassing approximately intention prompts for each of the most pertinent categories, namely marketing, covers and headers, posts, events, and advertising. In a complementary effort, we have compiled an extra set of around unique, creative user intentions. The graphic design images derived from these inventive prompts are displayed for your reference in last row of Figure 16. Refer to the supplementary for an all-inclusive review of these roughly intention prompts, along with the creative ones. Figure 9 furnishes comprehensive statistics of our benchmark from myriad perspectives. We conduct all the experiments on DesignerIntention by default.
4.2 Main Results
Comparison with State-of-the-art
Figure 15 presents a comparison with various state-of-the-art text-to-image generation models that are augmented with GPT-. The usage of GPT- stems from the fact that traditional text-to-image models generally necessitate detailed descriptions, which are often missing in the intention prompt. Consequently, we utilize GPT- across all models to translate the user’s simple intentions into intricate captions. The generation results are then evaluated accordingly across different systems. We provide the original intention prompts in Table 1. We provide a detailed enhanced prompts, augmented using GPT-, within the supplementary material.
In addition, we showcase their predicted numerical quality scores from GPT-4V(ision) based on quality assurance prompt in Figure 10. A careful evaluation reveals that our Cole achieves very competitive quality accessing results even compared to the latest DALLE3. Our Cole achieves the best quality when generating covers & headers or posters. An interesting observation is that the background image generated by DALLE3 is of exceptional quality but the text regions are non-editable. We also empirically show some impressive results of our Cole when replacing our text-to-background generation model with DALLE3 by leveraging the text segmentation models, inpainting models, and others.
Seed #1
Seed #2
Seed #3
Seed #4
Seed #5
initial image
+ object editing
+ text editing
+ combined editing
@@ 4.2 Main Results
Qualitative results
Figure 16 showcases a variety of graphic design images entirely created by our system. It’s noteworthy that our system performs exceptionally well, considering that each model within it is trained independently in a hierarchically organized manner. However, we recognize several limitations within our Cole system, including: (i) the arrangement of the typography block, (ii) the limited number of editable visual elements in the image, and (iii) the restricted diversity in typography color selection. Addressing these issues is a direction we’d like to pursue in our future work.
@@ 4.3 Ablation Experiments
Design LLM
We assess the performance of the Design LLM by calculating the semantic textual similarity (STS) between the predicted text and the ground-truth ones using Sentence-BERT [28].
This evaluation is done across various keys, such as global or background or object caption, keywords, and heading&sub-heading&bodytexts, following the generated JSON file. Their average STS scores are , , , , and , respectively.
Text-to-Background & Text-to-Object
Figure 12 showcases the generated background image layer, object image layer, and the composed image layer. The displayed visualizations highlight the proficiency of our text-to-object generation model in crafting a coherent object image layer and its corresponding alpha mask. We have also carried out a detailed comparison of the background images generated by our model and others [26, 19, 2]. Refer to the supplementary material for more details.
Typography LMM & Reflect LMM
We visualize the predicted results with the ground-truth images as the condition in Figure 12. We assess these results from two perspectives: typography scoring with GPT-4V(ision) and the localization accuracy including , , and . For a more comprehensive comparison and the ablation of the reflect-LMM, refer to the supplementary material.
Design Diversity
Designers often face the critical challenge of delivering diverse designs while maintaining high quality. Our Cole system empirically excels in this regard, as evidenced by its operation with various random seeds shown in Figure 14.
Flexible Editing
Our Cole offers a versatile and interactive platform that allows users to edit both the object image layer and typography. This functionality augments the capability to manipulate generated graphic design images from various perspectives. Figure 14 showcases examples of different editing operations. For instance, users have the flexibility to resize, reposition, or replace the object image layer, as well as to freely modify the font type and size.
@@ 5 Conclusion
This paper introduces an exceptionally efficient hierarchical generation framework., Cole, that simplifies the intricate process of graphic design generation by breaking it down into manageable, coordinated sub-tasks, each yielding distinct outputs. Leveraging the benefits of this hierarchical task decomposition, our system can generate competitive graphic design images from simple user intent even when compared to DALLE3 on the challenging DesignerIntention benchmark. The core strengths of our Cole system lie in ensuring precise text rendering and providing flexible editing options. We have also pinpointed several key limitations of our Cole system in the context of generating graphic design images, and we plan to tackle these in future research.
Acknowledgement We extend our gratitude to Xinyang Li for his invaluable assistance with data preparation. Our thanks also go to Yang Ou for providing insightful comments from a professional designer’s perspective. We appreciate the enlightening discussions offered by Chunyu Wang, Jianmin Bao, Mingxi Cheng, and Chong Luo. Additionally, we want to give a special mention to Weicong Liang for his indispensable help in addressing the LaTeX compilation errors.
@@
A Graphic Design Pipeline
The generation of high-quality graphic designs from user intentions requires the harmonious integration of several roles or components, each tailored to address specific aspects of the graphic design process [36, 34]. We initially offer a detailed plan that outlines these roles, clarifies their purposes, and explains their interactions. Figure 17 illustrates the complete pipeline.
Creative Design and Interpretation
@@ A Graphic Design Pipeline
The primary function of this role involves understanding user intentions and translating them into tangible design elements like titles, subtitles, headings, subheadings, and body text. It also sets the tone and style of the graphic design, which incorporates elements such as primary color, theme, style, and captions.
Visual Composition and Enhancement
@@
The key responsibility of this role is to uphold the principles of design aesthetics and layout, which involves the generation, assessment, and adjustment of visual appeal and spatial organization. The generation process mainly utilizes a comprehensive library of images, icons, and textures to enhance the design. This module further refines the initial concepts originating from the Creative Design and Interpretation engine, optimizing the layout and enriching the design with suitable assets. For instance, it is imperative to not only generate images of objects, decorations, and backgrounds, but also to predict various typographic attributes for elements such as the heading, subheading, and body text.
Feedback and Quality Assurance System
@@
The principal function of this role is to integrate user feedback into the design in an iterative manner, with the objective of ensuring that generated designs comply with quality benchmarks and regulatory standards. This module guarantees that the designs not only align with user expectations, but also fulfill professional and legal criteria. It maintains a continuous interaction with the first two modules to guide the design refinement process.
Collaboration Loop
@@
This system ensures that the designs not only meet user expectations, but also adhere to professional and legal standards. It maintains a constant dialogue with the first two modules to navigate the design refinement process effectively.
Discussion
@@
Unlike traditional graphic design systems that rely significantly on the expertise of professional designers, we aspire to create a DesignerAgent. This innovative system is designed to generate superior graphic designs, leveraging user intention descriptions alone. By fine-tuning state-of-the-art Large Multimodal Models (LMMs), Large Language Models (LLMs), and diffusion models, we aim to fully exploit our carefully curated, high-quality graphic design datasets to achieve this objective.
B. Detailed System Prompts for GPT-3.5 and GPT-4V(ision)
@@
We introduce the Intention Generation Prompt for GPT3.5, Quality Assurance Prompt for GPT-4V(ision), and Pair-wise Comparison Prompt for GPT-4V(ision) as follows: The Intention Generation Prompt is crafted to deduce the user’s purpose utilizing diverse image data, such as keywords, all text visible in the image, the image’s title, among others. The Quality Assurance Prompt for GPT-4V(ision) is established to evaluate the quality of the created graphic design images across roughly five dimensions. These dimensions include design and layout, content pertinence, graphics and imagery, creativity, and the use of typography and color. The Pair-wise Comparison Prompt for GPT-4V(ision) is proposed to objectively compare our generated images with results from other methods from six perspectives of graphic design.
C. The Detailed List of DesignerIntention

### 18 Mar 2024 v2

@@ 3.4 Text-to-Object Diffusion Model: Visual Reasoning based on the Generated Background Image
Preparing Text-Object-Background-AlphaMask Quadruples To train a text-to-object diffusion model capable of perceiving both the object captions present in the JSON file and the visual pixels predicted by the text-to-background diffusion models, we prepare the (object-caption, background-image, object-image, alpha-mask) quadruples. We extract the object-image and the alpha-mask from the raw data, while the object-caption and the background-image are prepared following the above-mentioned scheme.
@@
Implementation and Training Settings Following the instruct-P2P approach [3], we concatenate the noise latent vectors with the latent vectors of the conditioning background image with the alpha mask for the object layer, and we add additional input channels to the first convolutional layer. All available weights of the diffusion model are initialized from pre-trained checkpoints, while weights operating on the newly added input channels are initialized to zero. As a result, we can predict both the object image and its alpha mask layer, which enables us to control the transparency of the object pixels. Consequently, we can seamlessly blend the object image layer with the background image layer in accordance with the alpha mask. To ensure a more reasonable placement of object pixels and coherence in style with the background image, we propose predicting an additional composed image. This is achieved by adding three extra channels to both the input noise vectors and the output prediction, a step that proves critical for accurate object layer prediction. Therefore, our text-to-object diffusion model needs to predict channels, where the object image consists of channels, the object mask consists of channel, and the composed image consists of channels. Besides, we also increased the maximum number of text tokens from to following the text-to-background diffusion model.
@@
Similar to the text-to-background generation model, we optimize our assembled dataset, comprising approximately (object-caption, background-image, object-image, alpha-mask) quadruples. The relatively smaller number of quadruples is due to the frequent absence of the object image layer in many graphic design images. We employ the instruct-P2P scheme for optimization, which is based on the DeepFloyd [18] IF diffusion models, and this process is carried out over an estimated span of epochs.
Qualitative Examples Figure 6 provides a selection of representative examples that demonstrate our method of decomposing complex graphic design images into distinct layers of background and object images. It’s important to note that predicting the object alpha mask is essential for facilitating flexible editing operations on the object pixels.
3.5 Typography LMM: Layout Planning and Attribute Reasoning for Visual Text
@@
Typography [37] is the art and technique of arranging type to make written language legible, clear, and visually engaging for the reader. It therefore requires complex reasoning and planning capabilities. It encompasses the design of font style, appearance, and structure, with the goal of eliciting specific emotions and conveying certain messages. In essence, typography brings text to life. Contrary to recent efforts, such as [31], [13], and [10], which primarily employ diffusion models to generate novel letter-level or word-level semantic or artistic typographies, our focus is more directed towards the utilization of an LMM (Large Multimodal Model). This approach allows us to address the highly challenging task of image-level typography. This task encompasses selecting font types, establishing hierarchical arrangements, and adjusting leading (line spacing), tracking (character spacing), and kerning (space between specific characters) based on the composed images. Consequently, it demands a robust capability for layout planning and attribute reasoning from the Typography LMM.
Preparing Composed Images and Typography Data We construct nearly triplets of data, each consists of multiple texts, composed images of backgrounds and objects, and typography information. Figure 7 displays a representative example of a typography JSON file, illustrating that it comprises attributes, with the value of ‘text’ already predicted by the Design-LLM. We extract all these typography attributes directly from the raw data of each graphic design image.
@@
Implementation and Training Settings In our experiments, we elected to fine-tune the LLaVA--B model [24] using nearly triplets for epochs. It is evident that predicting visually appealing typography information for each text box, based on the input image, is a complex task. This is due to the requirement of predicting approximately attributes for each text box. To our knowledge, even the most recent effort [29] only manages to predict typography attributes, assuming the other attributes are fixed. To simplify the prediction process, we further categorize the continuous floating numbers into discrete bins as per the method outlined in [6] and we also utilize a unique token to represent the values that fall within the same bin. We explain more details of the bin partition scheme in the supplementary material.
3.6 Multi-Layered SVG Editor and Renderer: Support Layer-wise Flexible User Editing
After generating all the design image layers, we propose compiling them into a single SVG file using the open-source tool SVG-Edit[30]. Following this, we opt to utilize Justinmind222https://www.justinmind.com/, a tool that facilitates a wide range of flexible, layer-wise user editing operations. These operations include repositioning and resizing text boxes or object layers. Additionally, Justinmind allows for modifications to be made to font size, color, type, and other attributes within each text box, enhancing the user’s ability to customize their designs further. We present the individual image layers and the fully composed image in Figure 1. Additionally, we showcase a series of representative examples that demonstrate the outcomes following various layer-wise editing operations in Figure 9.
@@ 3.7 Reflect LMM & Quality LMM
Considering that initial multi-layered graphic design outputs may not always align with user expectations, we further propose the integration of an added Reflect-LMM. This component is designed to critically analyze the generated design images and predict necessary adjustments to the JSON file, thereby further improving the resulting graphic designs.
@@
Reflect LMM The entire self-reflection framework is illustrated on the left side of Figure 9. As an integral part of our implementation plan, we fine-tune a LLaVA--B model on a dataset that comprises more than artificially generated noisy JSON files. Each of these files is paired with a corresponding ground-truth JSON file. Our primary modification involves the addition of random shift noise to the position of each ground-truth typography block. The key insight lies in modeling the conditional transformation of typography attributes from suboptimal to ideal, leveraging the understanding of the rendered composite images.
Quality LMM We utilize GPT-4V(ision) [25] to evaluate the quality of the generated graphic design images, as displayed in the “Quality Assurance Prompt for GPT-4V(ision)” in the appendix. Through empirical analysis, we have found that GPT-4V(ision) excels at evaluating the quality of main aspects such as design and layout, content relevance, graphics and images, and innovation. However, it falls short in assessing the quality of typography and color. In addition, we also incorporate the evaluation prompt design as introduced in a recent concurrent effort [17]. This evaluates the graphic design images from approximately aspects, including text rendering quality, composition and layout, color harmony, cinematography, style, image-text alignment, aesthetics, and overall design.
4 Experiment
(a) covers & head.
@@ 4 Experiment
Seed1
Seed2
Seed3
Seed4
To assess our Cole’s capability for design-oriented planning and reasoning in generating graphic design images, we conducted extensive experiments on our proposed DesignerIntention, which consists of approximately graphic design intention prompts. Leveraging GPT-4V(ision), we conducted a multi-faceted comparison [2, 17] between our approach and the state-of-the-art image generators DeepFloyd/IF, SDXL, DALLE3, and CanvaGPT, which utilize GPT- as an intention interpreter. Subsequently, we assembled approximately 20 general users and 6 professional designers to conduct a subjective assessment using COLE, comparing it separately with both DALLE3 and CanvaGPT. Then we show that the zero-shot performance with our Typography-LMM even outperforms the latest layout planning methods in Crello text box placement task. Last, we carry out ablation experiments to scrutinize the performance of each component and discuss the limitations of our system.
4.1 DesignerIntention Benchmark
We have constructed a benchmark for designer intentions, amassing approximately intention prompts for each of the most pertinent categories, namely marketing, covers and headers, posts, events, and advertising. In a complementary effort, we have compiled an extra set of around unique, creative user intentions. Refer to the supplementary material for an all-inclusive review of these approximately intention prompts, along with the approximately creative ones, and comprehensive statistics. We conduct all the experiments on DesignerIntention by default.
4.2 Main Results
Comparison with State-of-the-art Figure 11 presents a comparison with various state-of-the-art text-to-image generation models that are augmented with GPT-. Notably, we also include the results of the widely recognized custom CanvaGPT [4], which integrates GPT-4 with Canva’s templates. Each Canva template fully produced by professional designers incurs significant costs. The usage of GPT- stems from the fact that both traditional text-to-image models and CanvaGPT generally necessitate detailed descriptions, which are often missing in the intention prompt. Consequently, we utilize GPT- across all models to translate the user’s simple intentions into intricate captions. The generation results are then evaluated accordingly across different systems. We provide the original intention prompts in Table 2. We provide a detailed enhanced prompts, augmented using GPT-, within the supplementary material.
In addition, we showcase their predicted numerical quality scores from GPT-4V(ision) based on quality assurance prompt in Figure 11. A careful evaluation reveals that our Cole achieves very competitive quality accessing results even compared to the latest DALLE3 and CanvaGPT. Our Cole secures leading performance across three critical dimensions: design layout, typography color, and innovation, consistently achieving second-best in quality. An interesting observation is that the background image generated by DALLE3 is of exceptional quality but the text regions are non-editable. Figure 1 shows some impressive results of our Cole when replacing our text-to-background generation model with DALLE3 by leveraging the text segmentation models, inpainting models, and others.
Qualitative results Figure 12 showcases a variety of graphic design images entirely created by our Cole system. It’s noteworthy that our system performs exceptionally well, considering that each model within it is trained independently in a hierarchically organized manner. However, we recognize several limitations within our Cole system, including: (i) the arrangement of the typography block, (ii) the limited number of editable visual elements in the image, and (iii) the restricted diversity in typography color selection. Addressing these issues is a direction we’d like to pursue in our future work.
@@ 4.3 Ablation Experiments
Text-to-Background & Text-to-Object Figure 15 illustrates the critical role of incorporating composed image prediction into the text-to-object diffusion model. This integration markedly enhances the model’s ability to plan layouts. Furthermore, we have conducted an exhaustive comparison between the background images produced by our model and those generated by other methods, such as those described in [26, 18, 2]. We found that most alternative approaches often result in objects that fill the entire rectangular canvas. For additional information, please consult the supplementary material.
Typography LMM & Reflect LMM We visualize the predicted results with the ground-truth images as the condition in Figure 14. We assess these results from two perspectives: typography scoring with GPT-4V(ision) and the localization accuracy including , , and . To assess the generalization capability of our approach in layout planning and attribute reasoning for visual text, we evaluate the zero-shot performance on the Crello benchmark and summarize the comparison results in Table 3. Our Typography-LMM outperforms the previous model, FlexDM, by +4.5 and +6.2 in terms of IoU score for single and multiple text box placement tasks, respectively Refer to the supplementary material for the ablation of reflect-LMM.
Design Diversity Designers frequently confront the challenge of producing varied yet high-quality designs. Our Cole system demonstrates exceptional performance in this aspect, as illustrated in Figure 14. We provide additional ablation experiments on the Design LLM and others in the supplementary material.
@@ 5 Conclusion
This paper introduces an exceptionally efficient hierarchical generation framework, Cole, that simplifies the intricate process of graphic design generation by breaking it down into manageable, coordinated sub-tasks, each yielding distinct outputs. By leveraging design-oriented visual planning, reasoning, and layer-wise generation capabilities during this hierarchical task decomposition, our system can generate multi-layered and editable graphic design images. We demonstrate the advantages of our approach by comparing it to the most powerful systems like DALLE3 and CanvaGPT. Additionally, we have pinpointed several key limitations of our Cole system in the context of generating graphic design images, and we plan to address these in future research.
Acknowledgement We extend our gratitude to Xinyang Li for his invaluable assistance with data preparation. Additionally, we are grateful to Yang Ou, along with both designers and non-designers, for their insightful feedback from a professional design and normal user perspective. We appreciate the enlightening discussions offered by Chunyu Wang, Jianmin Bao, Mingxi Cheng, and Chong Luo. Additionally, we want to give a special mention to Weicong Liang for his indispensable help in addressing the LaTeX compilation errors.
@@
A Graphic Design Pipeline of Human Designer
The creation of high-quality graphic designs from user intentions necessitates a seamless integration of diverse roles or components, each uniquely designed to tackle specific facets of the graphic design process [35, 33]. Initially, we present an elaborate blueprint that delineates these roles, elucidates their functions, and delineates their interactions. Figure 16 depicts this comprehensive pipeline, serving as the conceptual foundation for the design of our Cole system.
@@ A Graphic Design Pipeline of Human Designer
Creative Design and Interpretation The core responsibility of this role is to interpret user intentions and convert them into specific design elements, including titles, subtitles, headings, subheadings, and body text. This role also establishes the tone and style of the graphic design, encompassing aspects like primary color, theme, style, and captions.
@@
Visual Composition and Enhancement The key responsibility of this role is to uphold the principles of design aesthetics and layout, which involves the generation, assessment, and adjustment of visual appeal and spatial organization. The generation process mainly utilizes a comprehensive library of images, icons, and textures to enhance the design. This module further refines the initial concepts originating from the Creative Design and Interpretation engine, optimizing the layout and enriching the design with suitable assets. For instance, it is imperative to not only generate images of objects, decorations, and backgrounds, but also to predict various typographic attributes for elements such as the heading, subheading, and body text.
@@
Feedback and Quality Assurance System The principal function of this role is to integrate user feedback into the design in an iterative manner, with the objective of ensuring that generated designs comply with quality benchmarks and regulatory standards. This module guarantees that the designs not only align with user expectations, but also fulfill professional and legal criteria. It maintains a continuous interaction with the first two modules to guide the design refinement process.
@@
Collaboration Loop This system ensures that the designs not only meet user expectations, but also adhere to professional and legal standards. It maintains a constant dialogue with the first two modules to navigate the design refinement process effectively.
@@
Discussion Unlike traditional graphic design systems that rely significantly on the expertise of professional designers, we aspire to create a DesignerAgent. This innovative system is designed to generate superior graphic designs, leveraging user intention descriptions alone. By fine-tuning state-of-the-art Large Multimodal Models (LMMs), Large Language Models (LLMs), and diffusion models, we aim to fully exploit our carefully curated, high-quality graphic design datasets to achieve this objective.
B Detailed System Prompts for GPT-3.5 and GPT-4V(ision)
@@
We introduce the Intention Generation Prompt for GPT-3.5, Quality Assurance Prompt for GPT-4V(ision), and Pair-wise Comparison Prompt for GPT-4V(ision) as follows: The Intention Generation Prompt is crafted to deduce the user’s purpose utilizing diverse image data, such as keywords, all text visible in the image, the image’s title, among others. The Quality Assurance Prompt for GPT-4V(ision) is established to evaluate the quality of the created graphic design images across roughly five dimensions. These dimensions include design and layout, content pertinence, graphics and imagery, creativity, and the use of typography and color. The Pair-wise Comparison Prompt for GPT-4V(ision) is proposed to objectively compare our generated images with results from other methods from six perspectives of graphic design.
C The Detailed Prompt List of DesignerIntention